# Quantum Entanglement without nonlocal causation in (3,2)-dimensional spacetime

### v3

@@ Abstract
This work aims at exploring whether the nonlocal correlations due to quantum entanglement could exist without nonlocal causation. This is done with the aid of a toy model to investigate whether the ability of two quantum entangled particles to ”correlate” their behaviors even at very large distances and in the absence of any physical connection can be seen as due to an exchange of information through an extra-temporal dimension. Since superluminal information exchange is forbidden in our (3,1) space-time, an extra-temporal dimension is needed to recover the physical picture of finite velocity information exchange between entangled entities. Assuming that the geometry of space-time of dimension (3,2) is described by a metric containing a warping factor, the confinement of the massive particles in the extra time dimension follows. Therefore, why we do not experience an infinitely large extra time dimension can be explained. The toy model proposed here is defined by borrowing Bohm-Bub’s proposal to describe the wavefunction collapse using nonlinear (non-unitary) dynamical equations and then elaborating this approach for an entangled system. The model so obtained is just speculative without any claim of being robust against any criticism, nevertheless, it satisfies the purpose of giving the possibility to the hypotheses formulated above to be verified experimentally; in fact, it proposes an experiment potentially interesting which would otherwise be immediately dismissed as manifestly trivial. The proposed experiment would consist of checking the possible violation of Bell’s inequality between two identical but independent systems under appropriate conditions. Beyond its theoretical interest, entanglement is a key topic in quantum computing and quantum technologies, so any attempt to gain a deeper understanding of it could be useful.
@@ I Introduction
Einstein’s claim of incompleteness of quantum mechanics became susceptible of an experimental confirmation or refutation after John Bell’s proposal of a quantitative criterion bell to test the completeness of the theory versus the need of the so-called hidden variables to recover a locality condition fulfilling Einstein’s causality. Under a suitable locality assumption, Bell’s theorem states that local hidden-variables theories are constrained by some given inequality, and some predictions of quantum mechanics can violate this inequality. A pioneering experimental implementation of a test based on Bell’s criterion was put forward by Clauser, Horne, Shimony and Holt clauser and led to a confirmation of quantum mechanics through the violation of Bell’s inequality. However, having held fixed the orientation of the polarizers during the experimental run of this and other experiments made the results not conclusive because in the case of static experiments both the locality condition and the validity of Bell’s inequality can be questioned. Bell thus insisted on the opportunity of performing experiments according to the Bohm-Aharonov proposal bohm of changing the settings of the instruments while the correlated particles were in flight. In so doing, the locality condition is entailed by Einstein’s causality forbidding faster-than-light interactions.
Then a strict and undeniable violation of Bell’s inequality was given by Alain Aspect’s experiments aspect1 ; aspect2 where each single channel polarizer was replaced by a fast switching device redirecting the incident light beam to a differently oriented polarizer. The detection events on opposite sides of the experimental apparatus were separated by a space-like interval, thus free of locality loopholes. Further confirmations of the violation of Bell’s inequality were then obtained at increasingly large spatial separations nl1 ; nl2 ; nlz ; nl3 ; nl4 . Therefore, quantum nonlocality is undeniably a property of our physical world, but Bohr’s viewpoint can sound tautologic and unsatisfactory compared to our description of physical phenomena based on events located in space and evolving in time, so that the existence of correlations between events separated by an arbitrarily large space-like interval is a conundrum that goes against our intuitive perception of the physical world. Of course, not only quantum mechanics compels us to give up our intuitive perception of the physical world, nevertheless we can wonder whether - to some extent - we can recover a ”causal” description of quantum nonlocality by paying a different price with respect to Bohr’s claim of indivisibility of quantum systems.
The price to pay in the framework proposed in the present work is to consider the existence of an extra time dimension. In what follows there is no pretence of proposing something that could qualify as a new theory, rather, by means of a toy model we want to move some first steps in exploring whether nonlocality can be given a ”more familiar” interpretation by considering that while performing the measurement of an entangled state this appears as actually composed of independently existing units that exchange information via a hidden sub-quantum field ${\mathscr{X}}({\bf x},t,\tau)$ that propagates at finite velocity through an extra dimension of temporal kind $\tau$. All this while respecting quantum contextuality, the property of quantum phenomenology proved by the Bell-Kochen-Specker theorem BKS ; KS . This idea can be tested against an experiment that, if proved feasible and produced the hypothesized result, would make worthwhile and necessary to go beyond the toy model discussed below.
@@ I.1 A quick digression on extra time-dimensions
Summarizing, the possibility for a local dynamical theory in more dimensions than 3+1 to generate fundamentally non-local effects in lower dimensional space has been shown in several works where the extra dimensions (also additional times) play the role of hidden variables (see also Ref. physrepGenovese ).
@@ II.2 Extending Bohm-Bub’s model to entangled particles
according to the initial sign of the terms in the first parenthesis of the r.h.s. of each one of the equations above, it is immediately evident that if the first particle collapses to the state |+ket|$\rangle$| + ⟩ the second one collapses to |ket|$\rangle$| - ⟩ and viceversa. In analogy with the original Bohm-Bub’s model, the $\xi^{(j)}_{\pm}(t)$ are assumed to be hidden random variables, almost everywhere non vanishing, and constant during measurement. The non-locality is here expressed by the fact that even if the two particles are separated by a space-like interval, the values of the quantities $R^{(j)}$ have an instantaneous mutual influence on the respective wavefunction collapses. Hereafter our main hypothesis enters the game. We assume that the wavefunctions collapses of the two particles - that now have their own individuality - are correlated via an exchange of information mediated by a hidden sub-quantum field ${\mathscr{X}}({\bf x},t,\tau)$ which propagates through an extra temporal dimension described by the variable $\tau$. We are replacing the hidden variables $\xi^{(j)}_{\pm}(t)$ with a completely different physical entity, that is, a real hidden physical field ${\mathscr{X}}({\bf x},t,\tau)$, almost nowhere vanishing (that is with the possible exception of a zero measure subset of space), a condition that could be ensured for example by the existence of a stochastic background component of the field. We tentatively assume that the dynamical evolution of the scalar field ${\mathscr{X}}({\bf x},t,\tau)$ is described by a wave equation
@@
that is a wave equation where a tentative nonlinearity is introduced to account for the propagation of a non dispersive impulse, actually a sine-Gordon soliton, in a radial direction $r$ (under spherical symmetry), stemming from a given point ${\bf x}$, and carrying the information coded by the source term ${\widetilde{F}}[{\bf x},\psi_{\pm}^{(j)}(t),\tau;t_{0},\tau_{0}]$, where $t_{0}$ and $\tau_{0}$ stand for the initial times. In analogy with the standard assumption belinfante for the evolution of $\xi(t)$, that is, $\dot{\xi}(t)=0$ during an impulsive measurement, no term containing the derivative of ${\mathscr{X}}({\bf x},t,\tau)$ with respect to our usual time is considered in the equation above. At the present state of affairs, we are tackling a toy model because we lack physical input to make more definite formulations, in particular to assign the analytical form of the function ${\widetilde{F}}[{\bf x},\psi_{\pm}^{(j)}(t),\tau;t_{0},\tau_{0}]$, nevertheless, for the moment being it is sufficient to formally enter this term in Eq.(20) to represent the source of the information carried by ${\mathscr{X}}({\bf x},t,\tau)$. In fact, already at the present stage an experiment seems conceivable (see the next section) to test whether or not the surmised scenario corresponds to the physical reality, in case of a positive experimental outcome one could make more motivated assumptions. The collapse equations (17) are then rewritten as
@@
It is important to remark that the modification of the Schrödinger equation (1) is without any consequence on the standard unitary evolution of a quantum system - including the Bell state (14) - until $\gamma(t)=const=N_{0}\lambda_{GRW}$ takes a large value because of the interaction with the measuring apparatus. In other words, neither the field ${\mathscr{X}}({\bf x},t,\tau)$ nor the extra temporal dimension have any consequence on the standard unitary evolution until measurement.
@@
The final outcome of the measure is determined by the evolution of $|\Phi^{(1)}\rangle|\Phi^{(2)}\rangle$ which depends on the values that the field ${\mathscr{X}}({\bf x},t_{0},\tau_{0}+\Delta\tau)$ had immediately before the measurement at ${\bf x}_{1}$ and ${\bf x}_{2}$, and thus it depends on the random component of the field and on the unknown details about the shaping of the information-carrying impulses. All this makes the individual outcomes of measurements on entangled entities unpredictable, outcomes which however remain perfectly correlated. In other words, although it is demonstrated that the Bell inequality is violated by non-product states [43], nevertheless, also with the assumption done in the present work that a Bell state factorizes when both entangled entities interact simultaneously with the respective measurement devices, the physical information about the entanglement - and thus the violation of Bell inequality - is maintained by the collapse equations entailing correlated outcomes of the measures via the field ${\mathscr{X}}({\bf x},t,\tau)$.
@@ II.3 Proposal for a nonconventional experiment
The source term ${\widetilde{F}}[{\bf x},\psi_{\pm}^{(j)}(t),\tau;t_{0},\tau_{0}]$ in equation (20), even if of unspecified analytic form, amounts to assuming that the field ${\mathscr{X}}({\bf x},t,\tau)$ encodes the information about the wavefunction collapse due to the measurement performed on a given system, and equation (20) tells that this information is spread in every spatial direction. Therefore, a collapse information could also reach another identical system driving its collapse and mimicking entanglement correlation. This is the hypothesis to be experimentally proved or disproved that also suggests how to design an appropriate experiment. In fact, let us now consider two identical EPR sources, schematically depicted in Figure 3, simultaneously emitting each a pair of entangled photons or particles. Alice and Bob perform measurements at the ends of the first system, and Eve and Tom at the ends of the second system. Of course, the photons (or particles with spin) reaching Alice and Bob or Eve and Tom are in state superposition and constrained to fulfil a conservation law (momentum, energy, angular momentum) which is not the case of two particles reaching, say, Alice and Eve because a-priori the entities (say photons, as in Figure 3) reaching Alice and Eve are independent, that is, not entangled. Now, as soon as the photon detected by Alice has ”chosen” its polarization state how does the photon being detected by Bob ”know” that it has no choice left for its polarization state? The orthodox reply is that they are part of the same indivisible system even if they are very far apart, but if we assume that the information about the outcome of the measurement performed by Alice is broadcasted in every spatial direction (through the extra temporal dimension), a particle belonging to an identical system, like the one reaching Eve, could be driven to collapse into the same complementary state as in Bob’s case. Therefore, let us imagine an experimental setup where the measurement apparatuses of Alice and Eve are spatially close to one another and both of them much farther from the apparatuses of their companions Bob and Tom, so that the extra-time interval $\Delta_{1}\tau$ needed by the hypothetical field ${\mathscr{X}}({\bf x},t,\tau)$ to exchange information between Alice and Bob on one side, and Eve and Tom on the other side, is much longer than the interval $\Delta_{2}\tau$ needed to exchange information between Alice and Eve. Now, the global state vector of the two independent subsystems - Alice, Bob and Eve, Tom - is the product $|\Psi_{A,B}\rangle|\Psi_{E,T}\rangle$ where
@@
An important remark is in order. What is described in this Section aims to propose the conceptual scheme of a possible experimental verification of the physical hypothesis underlying this work on quantum entanglement. Indeed, the practical implementation of this proposal could make use of recent developments to design a more sophisticated experimental setup than the one sketched in Fig.3, resorting for example to the use of solid-state devices where the phenomenon of parametric down-conversion occurs inside a non-linear crystal generating entangled photons physrepGenovese ; mandel ; kwiat1 , as well as various methods to ensure loophole-free experiments physrepGenovese ; kwiat2 . For instance, a method for simultaneous generation of two independent pairs of entangled photons could be borrowed from Ref.zeilinger .

### v4

@@ Abstract
This work aims at exploring whether the nonlocal correlations due to quantum entanglement could exist without nonlocal causation. This is done with the aid of a toy model to investigate whether the ability of two quantum entangled particles to ”correlate” their behaviors even at very large distances and in the absence of any physical connection can be seen as due to an exchange of information through an extra-temporal dimension. Since superluminal information exchange is forbidden in our (3,1) space-time, an extra-temporal dimension is needed to recover the physical picture of finite velocity information exchange between entangled entities. Assuming that the geometry of space-time of dimension (3,2) is described by a metric containing a warping factor, the confinement of the massive particles in the extra time dimension follows. Therefore, why we do not experience an infinitely large extra time dimension can be explained. The toy model proposed here is defined by borrowing Bohm-Bub’s proposal to describe the wavefunction collapse using nonlinear (non-unitary) dynamical equations and then elaborating this approach for an entangled system. The model thus obtained aims to be a first step into unexplored territory, certainly a model that can be improved, but which already satisfies the purpose of giving the possibility to the hypotheses formulated above to be experimentally verified. The required experimental test has to resort to an unusual experiment which would otherwise be immediately dismissed as manifestly trivial. The proposed experiment would consist of checking the possible violation of Bell’s inequality between two identical but independent systems under appropriate conditions. Beyond its theoretical interest, entanglement is a key topic in quantum computing and quantum technologies, so any attempt to gain a deeper understanding of it could be useful.
@@ I Introduction
Einstein’s claim of incompleteness of quantum mechanics became susceptible of an experimental confirmation or refutation after John Bell’s proposal of a quantitative criterion bell to test the completeness of the theory versus the need of the so-called hidden variables to recover a locality condition fulfilling Einstein’s causality. Under a suitable locality assumption, Bell’s theorem states that local hidden-variables theories are constrained by some given inequality, and some predictions of quantum mechanics can violate this inequality. A pioneering experimental implementation of a test based on Bell’s criterion was put forward by Clauser, Horne, Shimony and Holt clauser ; freedman and led to a confirmation of quantum mechanics through the violation of Bell’s inequality. However, having held fixed the orientation of the polarizers during the experimental run of this and other experiments made the results not conclusive because in the case of static experiments both the locality condition and the validity of Bell’s inequality can be questioned. Bell thus insisted on the opportunity of performing experiments according to the Bohm-Aharonov proposal bohm0 of changing the settings of the instruments while the correlated particles were in flight. In so doing, the locality condition is entailed by Einstein’s causality forbidding faster-than-light interactions.
Then a strict and undeniable violation of Bell’s inequality was given by Alain Aspect’s experiments aspect1 ; aspect2 ; aspect3 where each single channel polarizer was replaced by a fast switching device redirecting the incident light beam to a differently oriented polarizer. The detection events on opposite sides of the experimental apparatus were separated by a space-like interval, thus free of locality loopholes. Further confirmations of the violation of Bell’s inequality were then obtained at increasingly large spatial separations nl1 ; nl2 ; nlz ; nl3 ; nl4 . Therefore, quantum nonlocality is undeniably a property of our physical world, but Bohr’s viewpoint can sound tautologic and unsatisfactory compared to our description of physical phenomena based on events located in space and evolving in time, so that the existence of correlations between spatially separated events is necessarily attributed to some mutual influence that propagates in space with a finite velocity. This physical description is completely lost with quantum entanglement that implies instantaneous correlations between events separated by an arbitrarily large distance, meaning that mutual influences propagate at infinite velocity. Since this is a conundrum that goes against our intuitive perception and description of the physical world, some authors suggested to recover a realistic local model of quantum mechanics by resorting to superluminal communication between entangled objects at finite velocity $v>c$ and under appropriate conditions bohm ; eberhard . A-priori this is not in contradiction with special Relativity because it can be shown that two observers performing experiments with entangled objects cannot use the corresponding quantum correlations to communicate at a faster-than-light speed bohm1 . This is referred to as no-signalling property of quantum correlations. One of the first such proposal was formulated surmising that the superluminal influences at a distance are exerted via mechanisms involving an ether and effects propagating in that ether eberhard . Said differently, the superluminal exchange of information leading to wave function collapse would be mediated by quantum tachyons, and to avoid causal paradoxes it has been shown that they have to propagate isotropically in a preferred reference frame with the same velocity in all directions liberati . The compatibility between superluminal propagation and the fundamental principles of causality and relativity have been investigated liberati and experimental attempts have been made to measure the velocity $v$ of quantum information in preferred reference frames scarani ; salart ; cocciaro yielding lower bounds for $v>c$. However, in Ref.bancal it was shown that ”any possible explanation of quantum correlations in terms of influences propagating at any finite speed $v$” with $c<v<\infty$ could violate the ”impossibility of using non-local correlations for superluminal communication”. Hence to avoid conflict with Relativity theory $v$ cannot be finite and superluminal. The authors of this work conclude by affirming ”[…] to keep no-signalling, […] quantum non-locality must necessarily relate discontinuously parts of the universe that are arbitrarily distant. This gives further weight to the idea that quantum correlations somehow arise from outside spacetime, in the sense that no story in space and time can describe how they occur”. Although logically unescapable, this conclusion according to which a physical phenomenon originates outside of spacetime is disorienting and a source of conceptual discomfort, therefore, we can ask ourselves whether the statement ”arise from outside [ordinary] spacetime” can be replaced by ”arise from inside an enlarged spacetime”, in other words - after the ”no go” for superluminal quantum communication at finite velocity - we can try to recover a ”causal” description of quantum nonlocality by assuming the existence of an extra time dimension. It is important to remark that we are not proposing a method that evades the requirement for a superluminal exchange of information, the projection to the ordinary (3+1) spacetime of the information exchange between two entangled objects through an extra temporal dimension of (3+2) spacetime is superluminal but occurring at infinite velocity, thus respecting the no-signalling condition without any need for preferred frames. That is, the idea put forward in the present work leaves unchanged the nonlocality experienced in the ordinary (3+1) spacetime. In what follows there is no pretence of proposing something that could qualify as a new theory, but rather to put forward an unprecedented thought-provoking proposal to delve deeper into the mystery of quantum entanglement without contradicting the current explanation of quantum mechanics in any aspect. From Dirac’s axiomatic formulation of quantum mechanics in 1930 to nowadays quantum entanglement is considered a primitive (i.e. ”de facto”) property of quantum formalism Qaxiom . Aside from the failed proposal of tachyon communication, no attempt has been made to explain entanglement from a more fundamental level. Therefore the proposal put forward in the present work is not an alternative explanation to something already existing. Actually, by means of a simple (toy) model we move some first steps in an uncharted territory by exploring whether nonlocality can be given a ”more familiar” interpretation by considering that while performing the measurement of an entangled state this appears as actually composed of independently existing units that exchange information via a hidden sub-quantum field ${\mathscr{X}}({\bf x},t,\tau)$ that propagates at finite velocity through an extra dimension of temporal kind $\tau$. All this while respecting quantum contextuality, the property of quantum phenomenology proved by the Bell-Kochen-Specker theorem BKS ; KS . This idea can be tested against an experiment that, if proved feasible and if it produced the hypothesized result, would make worthwhile and necessary to go beyond the toy model discussed below, and would pave the way to the development of a topic of fundamental theoretical importance that nowadays may also have practical implications for cutting-edge research into quantum technologies. For example, the reply to questions like “is the degree of entanglement weakening at increasing distance?” is ”no”, according to a formalism describing as an indivisible system two entangled objects even if sitting at opposite borders of our galaxy nota . However, this formalism was developed to describe microscopic systems, and wondering whether it can be somewhat upgraded is of conceptual interest and potentially practical relevance Qinternet .
@@ I.1 A quick digression on extra time-dimensions
Summarizing, the possibility for a local dynamical theory in more dimensions than 3+1 to generate fundamentally non-local effects in lower dimensional space has been shown in several works where the extra dimensions (also additional times) play a key role (see also Ref. physrepGenovese ).
@@ II.2 Extending Bohm-Bub’s model to entangled particles
according to the initial sign of the terms in the first parenthesis of the r.h.s. of each one of the equations above, it is immediately evident that if the first particle collapses to the state |+ket|$\rangle$| + ⟩ the second one collapses to |ket|$\rangle$| - ⟩ and viceversa. In analogy with the original Bohm-Bub’s model, the $\xi^{(j)}_{\pm}(t)$ are assumed to be hidden random variables, almost everywhere non vanishing, and constant during measurement. The non-locality is here expressed by the fact that even if the two particles are separated by a space-like interval, the values of the quantities $R^{(j)}$ have an instantaneous mutual influence on the respective wavefunction collapses. Hereafter we introduce the following working hypothesis which is the theoretical core that is being considered. We assume that the correlated wavefunctions-collapses of the two particles - that now have their own individuality - are driven by an exchange of information mediated by a hidden sub-quantum field ${\mathscr{X}}({\bf x},t,\tau)$ which propagates through an extra temporal dimension described by the variable $\tau$. This hypothesis necessarily requires the extra time dimension because a superluminal exchange of information is forbidden in our standard space-time. We are replacing the hidden variables $\xi^{(j)}_{\pm}(t)$ with a completely different physical entity, that is, a real hidden physical field ${\mathscr{X}}({\bf x},t,\tau)$, almost nowhere vanishing (that is with the possible exception of a zero measure subset of space), a condition that could be ensured for example by the existence of a stochastic background component of the field. Let us remark that the variables $\xi_{i}(t)$ of the original Bohm-Bub model, as well as the field ${\mathscr{X}}({\bf x},t,\tau)$ introduced here, play a role only during the nonlinear, non-unitary dynamics of the state reduction, thus they do not belong to the ”family” of hidden variables theories where quantum mechanics is assumed to be a statistical approximation of some unknown deterministic theory where the values of the observables are already defined and fixed by unknown, hidden variables.
We tentatively assume that the dynamical evolution of the scalar field ${\mathscr{X}}({\bf x},t,\tau)$ is described by a wave equation
@@
that is a wave equation where a tentative nonlinearity is introduced to account for the propagation of a non dispersive impulse, actually a sine-Gordon soliton, in a radial direction $r$ (under spherical symmetry), stemming from a given point ${\bf x}$, and carrying the information coded by the source term ${\widetilde{F}}[{\bf x},\psi_{\pm}^{(j)}(t),\tau;t_{0},\tau_{0}]$, where $t_{0}$ and $\tau_{0}$ stand for the initial times. In analogy with the standard assumption belinfante for the evolution of $\xi(t)$, that is, $\dot{\xi}(t)=0$ during an impulsive measurement, no term containing the derivative of ${\mathscr{X}}({\bf x},t,\tau)$ with respect to our usual time is considered in the equation above. At the present state of affairs, we are tackling a toy model because we lack physical input to make more definite formulations, in particular to assign the analytical form of the function ${\widetilde{F}}[{\bf x},\psi_{\pm}^{(j)}(t),\tau;t_{0},\tau_{0}]$, nevertheless, for the moment being it is sufficient to formally enter this term in Eq.(20) to represent the source of the information carried by ${\mathscr{X}}({\bf x},t,\tau)$. In fact, already at the present stage an experiment seems conceivable (see the next section) to test whether or not the surmised scenario corresponds to the physical reality, in case of a positive experimental outcome one could make more motivated assumptions.
Let us now rewrite Eq.(1) as
$\frac{\partial\Psi({\bf x},t)}{\partial t}={\widetilde{\mathscr{B}}}({\bf x},t,{\overline{\tau}})-\frac{i}{\hbar}\hat{H}\Psi({\bf x},t)$ (21)
and rewrite the non-unitary evolution term (2) as
~${\widetilde{\mathscr{B}}}({\bf x},t,{\overline{\tau}})=\gamma\sum_{i}\psi_{i}(t)\phi_{i}({\bf x})\sum_{j}|\psi_{j}(t)|^{2}({\widetilde{R}}_{i}-{\widetilde{R}}_{j})$ (22)
where the hypothetical information-field ${\mathscr{X}}({\bf x},t,\tau)$ enters the terms ${\widetilde{R}}$ by replacing the hidden variables $\xi_{i}(t)$. We remark that the term ~${\widetilde{\mathscr{B}}}({\bf x},t,{\overline{\tau}})$ is not continuously dependent on $\tau$ but it depends just on point values of ${\mathscr{X}}({\bf x},t,\tau)$ at some fixed extra time ${\overline{\tau}}=\tau_{0}+\Delta\tau$, as is specified in the following, therefore this explains why in Eq.(21) $\Psi({\bf x},t)$ is not replaced by $\Psi({\bf x},t,\tau)$, in other words this means that we do not need a two-time extension of standard quantum mechanics.
Then, with the notations of Eq.(22), the collapse equations (17) are replaced by
@@
Notice that the field ${\mathscr{X}}({\bf x},t,\tau)$ is assumed to propagate at finite (unknown) velocity $w$ according to equation (20) through the extra time $\tau$ only, and it is assumed to propagate at infinite velocity in our standard (3+1) spacetime. In fact, the denominators $|{\mathscr{X}}({\bf x}_{1,2}-{\bf w}\Delta\tau,t_{0},\tau_{0}+\Delta\tau)|^{2}$ entering the above defined functions ${\widetilde{R}}^{(1)}_{\pm}$ and ${\widetilde{R}}^{(2)}_{\pm}$ account for the propagation of ${\mathscr{X}}({\bf x},t,\tau)$ only in the extra time dimension because the usual time enters solely with $t_{0}$, the instant of measurement, thus implicitly assuming nonlocal instantaneous correlation in $t$ with the absence of a finite propagation speed. This is an important point to ensure the no-signalling condition discussed in the Introduction.
It is important to remark that the modifications of the Schrödinger equation in (1) and (21) is without any consequence on the standard unitary evolution of a quantum system - including the Bell state (14) - until $\gamma(t)=const=N_{0}\lambda_{GRW}$ takes a large value because of the interaction with the measuring apparatus. In other words, neither the field ${\mathscr{X}}({\bf x},t,\tau)$ nor the extra temporal dimension have any consequence on the standard unitary evolution until measurement.
@@
The final outcome of the measure is determined by the evolution of $|\Phi^{(1)}\rangle|\Phi^{(2)}\rangle$ which depends on the values that the field ${\mathscr{X}}({\bf x},t_{0},\tau_{0}+\Delta\tau)$ had immediately before the measurement at ${\bf x}_{1}$ and ${\bf x}_{2}$, and thus it depends on the random component of the field and on the unknown details about the shaping of the information-carrying impulses. All this makes the individual outcomes of measurements on entangled entities unpredictable, outcomes which however remain perfectly correlated. In other words, although it is demonstrated that the Bell inequality is violated by non-product states gisin , nevertheless, also with the assumption done in the present work that a Bell state factorizes when both entangled entities interact simultaneously with the respective measurement devices, the physical information about the entanglement - and thus the violation of Bell inequality - is maintained by the collapse equations entailing correlated outcomes of the measures via the field ${\mathscr{X}}({\bf x},t,\tau)$.
@@ II.3 Proposal for a nonconventional experiment
The source term ${\widetilde{F}}[{\bf x},\psi_{\pm}^{(j)}(t),\tau;t_{0},\tau_{0}]$ in equation (20), even if of unspecified analytic form, amounts to assuming that the field ${\mathscr{X}}({\bf x},t,\tau)$ carries the information about the wavefunction collapse due to the measurement performed at some given place, and equation (20) tells that this information is spread in every spatial directions through the extra time dimension $\tau$. Let us repeat once more that entanglement cannot be explained through an instantaneous information transfer in our (3,1) space-time. Then equations (II.2) describe the entanglement between two objects as due to their correlated collapses of the respective wavefunctions driven by a mutual transmission of information by means of the field ${\mathscr{X}}({\bf x},t,\tau)$ through the extra time $\tau$. Therefore, the collapse information of a given system could also reach another identical system driving its wavefunction collapse, so mimicking entanglement correlation, even if these two systems are not entangled according to the quantum formalism nota2 . Hence the violation of Bell inequality could be unexpectedly found also in apparently trivial situations. This is the hypothesis to be experimentally proved or disproved that also suggests how to design an appropriate experiment. An experiment that would appear meaningless if compared to the standard ways of considering the violation of Bell inequality reported in the literature. To the contrary, it will turn out meaningful if the working hypothesis advanced in the present paper has a counterpart in the physical world. In fact, let us now consider two identical EPR sources, schematically depicted in Figure 3, simultaneously emitting each a pair of entangled photons or particles. An experimental method for simultaneous generation of two independent pairs of entangled photons for example is reported in Ref.zeilinger . Alice and Bob perform measurements at the ends of the first system, and Eve and Tom at the ends of the second system. Of course, the photons (or particles with spin) reaching Alice and Bob or Eve and Tom are in state superposition and constrained to fulfil a conservation law (momentum, energy, angular momentum) which is not the case of two particles reaching, say, Alice and Eve because a-priori the entities (say photons, as in Figure 3) reaching Alice and Eve are independent, that is, not entangled. Now, as soon as the photon detected by Alice has ”chosen” its polarization state how does the photon being detected by Bob ”know” that it has no choice left for its polarization state? The orthodox reply is that they are part of the same indivisible system even if they are very far apart, but if we assume that the information about the outcome of the measurement performed by Alice is broadcasted in every spatial direction (through the extra temporal dimension), a particle belonging to an identical system, like the one reaching Eve, could be driven to collapse into the same complementary state as in Bob’s case.
We could wonder why should we assume that Eve’s particle reacts to the information broadcast from Alice’s particle, or vice versa, and why it would not be more natural to assume that Alice’s and Bob’s particles only respond to each other’s information exchange, and similarly for the pair Eve and Tom. After the basic assumption that Alice’s and Bob’s particles respond to each other’s information exchange via the field ${\mathscr{X}}({\bf x},t,\tau)$, which is transmitted by each particle when hitting its measuring device, when we move on to consider four identical measuring devices receiving identical particles emitted synchronously from identical sources, in the absence of any hypothesis about how the field ${\mathscr{X}}({\bf x},t,\tau)$ might encode from which particular observer the information about the collapse of the wave function comes, the simplest assumption is that Alice’s and Eve’s particles can also respond to each other’s information exchange via the field ${\mathscr{X}}({\bf x},t,\tau)$ under suitable conditions as follows. So, let’s imagine an experimental setup in which Alice’s and Eve’s measuring apparatuses are spatially close to each other and both much further away from the apparatuses of their companions Bob and Tom, so that the extra-time interval $\Delta_{1}\tau$ needed by the hypothetical field ${\mathscr{X}}({\bf x},t,\tau)$ to exchange information between Alice and Bob on one side, and Eve and Tom on the other side, is much longer than the interval $\Delta_{2}\tau$ needed to exchange information between Alice and Eve. Now, the global state vector of the two independent subsystems - Alice, Bob and Eve, Tom - is the product $|\Psi_{A,B}\rangle|\Psi_{E,T}\rangle$ where
@@
An important remark is in order. What is described in this Section aims to propose the conceptual scheme of a possible experimental verification of the physical hypothesis underlying this work on quantum entanglement. Indeed, the practical implementation of this proposal could make use of recent developments to design a more sophisticated experimental setup than the one sketched in Fig.3, resorting for example to the use of solid-state devices where the phenomenon of parametric down-conversion occurs inside a non-linear crystal generating entangled photons physrepGenovese ; mandel ; kwiat1 , as well as various methods to ensure loophole-free experiments physrepGenovese ; kwiat2 .